# Data Watermarking for Sequential Recommender Systems

### v1

@@ 3.2.2. Watermark Construction
Now we explain how we construct the watermark. A detailed description is shown in Algorithm 1. In the sequential recommendation, user’s recent interactions usually contribute more to the prediction result. To validate this assumption, we conducted an experiment with SASRec (Kang and McAuley, 2018) model on the MovieLens-1M dataset to check how previous items contribute to the embedding of a target item. For each item in a sequence, we pick all items before it, and compute their attention scores w.r.t. the target item at the last transformer layer. Then we average the attention scores according to the relative positions and present the result in Fig. 2. We can observe that, the embedding of an item largely depends on the items close to it. Therefore, if we want the model to memorize the watermark, we need to define the watermark $S_{wm}$ as a continuous, uninterrupted sequence of selected items. This approach forces the model to learn a strong dependency among watermark items, facilitating the memorization of the watermark.
As for the selection of the watermark items, we compute the popularity of each item by their number of interactions, and randomly select $l$ items from the top-10% unpopular items [^1] We empirically found that 10% is a good choice to achieve both high-level unnoticeability and discriminability.. The selected items are randomly ordered in $S_{wm}$. We next explain why we choose not to use popular items as watermark items. On the one hand, since the long-tail distributions of real-world recommendation datasets introduce the popularity bias to recommender systems, where popular items can be observed in most users’ recommendation lists, selecting popular items as the watermark items can raise the watermark success rate in oracle models. This reduces the discriminability of the watermark. On the other hand, popular items have highly informative and robust embeddings. Since the watermark is a sequence that does not follow the true data distribution, forcing popular items to establish correlations with the out-of-distribution watermark sequence is very difficult and may even bring unpredictable damage to the model performance, thus harming the unnoticeability of the watermark.
The watermark length $l$ should be as small as possible to enhance the unnoticeability. A long watermark can be easily detected by statistic methods such as frequent pattern mining (Han et al., 2001; Ayres et al., 2002). Empirically, we found that $l=3$ yields a very good performance in most cases, achieving a good watermark success rate as well as being able to bypass frequent pattern mining. We empirically found that, with a reasonable number of users (1% by default, introduced in Sec. 3.2.3) being embedded with the watermark, the frequency of the watermark subsequence $S_{wm}$ is smaller than the most frequent benign subsequence. For example, for MovieLens-1M dataset, $S_{wm}$ with length $l$=3 appears for 60 times, whereas the most frequent benign subsequence with length 3 appears for 121 times.
3.2.3. Candidate Users & Inserting Positions
We call those users for whom we embed the watermark into the sequence of items as candidate users. The number of candidate users is determined by a hyperparameter $p$ called inserting ratio, which denotes the percentage of candidate users over all users. For example, if $p$=1and the number of users $|\mathcal{U}|$=1,000, then there would be 10 candidate users. The hyperparameter $p$ should also be as small as possible to enhance unnoticeability. We will discuss the influence of $p$ in the experiment section. To ensure universality, we select candidate users randomly, so that the watermark would have less bias towards any particular user group, reducing the correlation between the watermark and the specific user behaviors, thus increasing the universal watermark success rate for arbitrary queries. We note that some datasets do not allow repeated items in one user sequence, i.e., an item would appear at most once in a sequence. In this case, we would only select users that do not contain the watermark items as candidate users. Since the watermark items are unpopular items, the number of users that have interacted with the watermark items would be very small. Therefore, we still have a large enough user pool to pick the number of candidate users given by $p$. For each candidate user, we insert $S_{wm}$ at a random position, which helps reduce the correlation between the watermark and any particular inserting position, and therefore contributing to the universality.
With the randomness introduced in candidate user selection and inserting position selection, the watermark would have less interference with the original clean sequences, thus alleviating the difficulty for the model to memorize the watermark, while preserving the model utility on the recommendation task.
@@ 3.2.4. Watermark Evaluation
When the dataset owner suspects that a target model is trained on their watermarked dataset, they can claim the ownership by querying the model with the watermark. Specifically, we assume that the dataset owner only has black-box access to the target model, i.e., they can only query the model and observe the output. Thus, to evaluate the watermark, the model owner can randomly select some benign user sequences and append the first $l-1$ watermark items $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$ to the end of those benign users, and then query the target model with them. The goal is to check the ranking of the last watermark item $i^{wm}_{l}$ in the output ranking lists. If $i^{wm}_{l}$ is ranked at the top, the model owner can safely claim the ownership.
@@ 4. User Watermarking
So far, we have explored the dataset watermarking which is used to protect the copyright of the entire dataset. We are also interested in the problem of user watermarking, where an individual user may want to protect his/her own data by embedding a watermark in it. This is a very challenging problem. The quality of the watermark can be largely influenced by multiple factors such as the interaction history of the target user, the distribution of the entire dataset, as well as the implementation of the model. As a pioneer work, we define the problem of user watermark and present our initial solution to it. More powerful user watermarking techniques are a promising future direction.
@@ 4.1. Problem Definition
We define the problem of user watermarking as to inject a watermark into a target user’s interaction sequence, so that a model trained on the dataset that involves this watermarked user would have a high watermark success rate on that watermark. We assume that the target user does not have access to other users’ interactions, i.e., the target user cannot view or edit other users’ interactions. The target user is only capable of determining his or her own interaction sequence by interacting with any item at any predetermined time slot. We also assume that the target user does not have any knowledge regarding the recommender system. However, we assume that the target user has access to the item popularity, e.g., how many times an item has been interacted by all users. This is a reasonable assumption on most e-commerce and social platforms (e.g., a user can see how many times an item has been purchased on most of the e-commerce platforms). The target user can leverage such information to help design the watermark.
The user watermark should also meet the requirements of universality, unnoticeability, and discriminability as mentioned in Sec. 3.1. But things become more challenging as we are only permitted to inject the watermark into one sequence with limited prior knowledge. Next, we present our analysis and solution to this problem.
@@
12:  Uniformly sample $l$ items from the most unpopular items
@@
4.2. DWRS-U
4.2.1. Overview
We define the user watermark in a similar way as the dataset watermark. The user watermark is a sequence of watermark items $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l}\}$ with length $l$. $S_{wm}$ would only be inserted into the target user’s interaction sequence. To facilitate the memorization of the watermark, the watermark items are sampled from those with the least number of interactions, and the inserting position is chosen as the position right before the most unpopular subsequence of the target user interaction sequence.
4.2.2. Watermark Item Selection
Since the watermark signal strength is largely reduced compared with dataset watermark as it only appears once in the entire dataset, we need to compensate for the strength loss by other means, e.g., choices of watermark items, the inserting position, as well as the watermark length.
A detailed description of our DWRS-U algorithm for user watermarking is shown in Algorithm 2. We still randomly select unpopular items as the watermark items. The rationale is basically similar with that in Sec. 3.2.2. Since the watermark only appears once in the entire dataset, we need the model to memorize the watermark at first sight. This requires the model to overfit the watermark. To achieve this, watermark items should appear less frequently in other sequences, thereby introducing less bias in the memorization of the watermark. Consider the case where we try to use a popular item in the watermark. Since the popular item appears frequently in other sequences, its embedding would receive a large amount of training signals and becomes robust and steady. It would be very difficult for the item to be used in the single watermark sequence and establish strong correlations with other watermark items. Therefore, only unpopular items are suitable to fill in the watermark. To further ensure the watermark strength, instead of using the top-10% unpopular items, we here select those items with the least number of interactions as the candidate watermark items. For example, if the most unpopular item has three interactions, then all those items with three interactions are considered as candidate watermark items. If the number of items with three interactions are smaller than the watermark length $l$, we can expand the candidate item pool to include those items with four interactions. This is approximately equivalent to using a much smaller percentage (top-1% or even top-0.1%, considering the specific number of all items and the watermark length $l$) rather than the top-10% to determine the candidate pool of the watermark items. Since users have access to the item popularity, this strategy is feasible to implement. In this way, the unpopularity of the watermark items would be further enhanced, thus improving the memorization as well as the discriminability of the watermark.
4.2.3. Watermark Length & Inserting Position
@@
For the watermark length $l$, a larger $l$ naturally enhances memorization, as it provides more dependencies suggested by Fig. 2. This would make the prediction of $i^{wm}_{l}$ easier to overfit given $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$. In fact, we have empirically verified that, under the setting of Fig. 2, by setting $l>30$ and using unpopular items as watermark items, the watermark is almost guaranteed to be successful with Recall@10 $>$ 0.9, no matter where the inserting position is. It suggests that we can always create a strong watermark by increasing the watermark length $l$. However, a long watermark would fail the requirement of unnoticeability and may increase the risk of being detected. Therefore, we next discuss the possibility of choosing a dedicated inserting position which can help alleviate the need for a long watermark.
@@
If we want to reduce the watermark length $l$, we need to compensate for the dependency loss by other means. Since we are evaluating the watermark by predicting $i^{wm}_{l}$ given $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$, let us now focus on the last watermark item $i^{wm}_{l}$. Fig. 2 suggests that items before $i^{wm}_{l}$ contribute significantly to its embedding (since their distances from $i^{wm}_{l}$ are short). However, note that $i^{wm}_{l}$ also contributes to the embeddings of the items after it. For convenience, we name a sequence of those items after $i^{wm}_{l}$ as $\mathcal{C}_{n}$, where $n$ denotes the number of items considered. Since the embeddings of items in $\mathcal{C}_{n}$ are obtained by aggregating information from previous items, learning to predict items in $\mathcal{C}_{n}$ in the training process would require updating the embedding of $i^{wm}_{l}$ as well. Therefore, $i^{wm}_{l}$ and $\mathcal{C}_{n}$ would also establish strong correlations, and we may compensate for the dependency loss of reducing $l$ by utilizing $\mathcal{C}_{n}$ to increase the dependency capacity. To achieve this, items in $\mathcal{C}_{n}$ should also be unpopular, as their embeddings are less robust and can be easily fused with the watermark information, thereby increasing the dependency capacity of $S_{wm}$. To this end, we propose to insert the watermark into a position right before the most unpopular subsequence in the target user sequence. For example, if given $n$=10, we would first iterate through the target user sequence and extract all subsequences with length 10. Then for each subsequence, we compute the average popularity (number of interactions) of items in it, and select the most unpopular subsequence as $\mathcal{C}_{n}$. The watermark would be inserted right before $\mathcal{C}_{n}$. In this way, we may shorten the length $l$ of the watermark $S_{wm}$. As $l$ decreases, the dependency capacity of $S_{wm}$ also decreases. Thus the model would try to compensate for the capacity loss by utilizing items in $\mathcal{C}_{n}$. When evaluating the watermark, we would discard $\mathcal{C}_{n}$ and append $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$ to the query sequence and evaluate the prediction of $i^{wm}_{l}$.
@@ 4.2.4. Discussions
Some considerations may be raised regarding the above method. On the one hand, it seems that we may manually insert some unpopular items after $i^{wm}_{l}$ to serve as $\mathcal{C}_{n}$. However, we argue that this does not achieve the goal of reducing the watermark length. Although $S_{wm}$ itself is indeed shortened, we still have to insert $n$ more items to form $\mathcal{C}_{n}$, and the total number of inserted items would become $l+n$, which is more than $l$. On the other hand, it seems that inserting $S_{wm}$ after the most unpopular subsequence can achieve the same effect of increasing the dependency capacity. But putting unpopular item before the watermark is actually equivalent to expanding the watermark with those items. One still has to include those items in the query when evaluating the watermark in order to achieve a high watermark validity. We are also aware of a weakness of this method. Specifically, if the target user does not have a subsequence filled with unpopular items, or if those items are not unpopular enough, this method may fail and we may have to increase the length of the watermark.
In addition, we note that we can use DWRS-U to do dataset watermarking. But as we will see in our experiment results, DWRS-D already achieves a very good watermarking quality (high watermark validity and model utility) with a short watermark length and a small inserting ratio. In this case, using DWRS-U to do dataset watermarking would not bring a significant improvement on the watermark quality, but would instead increase the chance of exposure, because the sampling strategies in DWRS-U can introduce potential bias that could be detected. Therefore, DWRS-D is generally preferred for dataset watermarking.
Table 1. Dataset Statistics

### v2

@@ 3.2.2. Watermark Construction
In this section, we discuss the detailed design of DWRS-D.
Inserting the Watermark
According to Fig. 3, the RF is determined by both the dataset and the model. As a defender, we should assume no knowledge about the model architecture that an attacker would use, thus it is impossible for us to obtain the true RF.
Now let us discuss how the true RF can affect the memorization of the watermark. Given a model and a dataset, suppose the true RF contains $r_{1}$ items before the target item, as well as $r_{2}$ items after the target item. The watermark length is denoted by $l$, so the length of the watermark body is $l-1$. One possible case is that the watermark covers the entire RF, for example, $r_{1}\leq l-1$ and $r_{2}=0$. In this case, all items that contribute to $y_{wm}$ are from $x_{wm}$. In this case, the watermark can be easily memorized. In other cases, for example $r_{1}>l-1$ or $r_{2}>0$, the RF contains filler items other than $x_{wm}$. They will contribute to the aggregation of $y_{wm}$, interfering with the watermark memorization. For example, a certain filler item might appear multiple times in different sequences, so the model would learn correlations between this item and $y_{wm}$, which would introduce additional bias and lead to performance degradation when evaluating the watermark using only $x_{wm}$. To alleviate this issue, we could force the filler items to follow a uniqueness constraint, that is, an item would appear at most once as a filler item. The complete algorithm is shown in Algorithm 1.
Specifically, we will first define a proper RF range of the specific dataset according to the attention heatmaps. For example, we can define the RF of ML-1M as five items before the target item and five items after the target item. We start with a random sequence and insert $S_{wm}$ into a random position. We record the filler items that are within the RF but do not belong to $S_{wm}$. Then in each iteration, we draw one sequence from other unselected sequences. For this sequence, we iteratively go through all positions, until we find a position such that the RF does not contain any filler item that has been recorded. We stop until the number of sequences reaches the predefined insert ratio $p$ or we cannot find any of the remaining sequences that meet the uniqueness constraint. In this way, there will be no duplicate filler items, so the noise introduced by the filler items will be alleviated.
Watermark Items
The watermark items in $S_{wm}$ can be any item of the defender’s choice. However, in general, unpopular items with fewer interactions are more favorable. On the one hand, unpopular items are less observed in benign users’ recommendation lists than popular items (Yu et al., 2023), which helps improve the discriminability of the watermark by achieving a low watermark success rate on oracle models. On the other hand, unpopular items receive few training signals due to the limited number of interactions they have. Therefore, they can be easily adapted to the watermark pattern. We will address this problem again in Sec. 5.4.1 and show experiment results in Appendix A for different popularity.
Watermark Length
The watermark length $l$ can also be any positive integer. But according to Fig. 3, only items in the RF can contribute to the target item, so increasing the length of the watermark would have limited performance gain on the watermark success rate if the watermark items fall out of the RF of $y_{wm}$. In addition, a long watermark would be easily detected by statistic methods such as frequent pattern mining (Han et al., 2001; Ayres et al., 2002). Therefore, a short watermark is more favorable in achieving high watermark utility and good unnoticeability.
@@ 3.2.3. Watermark Evaluation
When the dataset owner suspects that a target model is trained on their watermarked dataset, they can claim the ownership by querying the model with the watermark. Specifically, we assume that the dataset owner only has black-box access to the target model, i.e., they can only query the model and observe the output. Thus, to evaluate the watermark, the model owner can randomly select some benign user sequences and append $x_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$ to the end of those benign sequences, and then query the target model. The goal is to check the ranking of the response item $y_{wm}=i^{wm}_{l}$ in the output ranking lists. If $i^{wm}_{l}$ is ranked at the top, the model owner can safely claim the ownership.
@@ 4. User Watermarking
So far, we have explored the dataset watermarking to protect the copyright of the entire dataset. We are also interested in the problem of user watermarking, where an individual user may want to protect his/her own data by embedding a watermark in it. This is a very challenging problem. The quality of the watermark can be influenced by multiple factors such as the interaction history of the target user, the distribution of the entire dataset, as well as the implementation of the model. As a pioneer work, we define the problem of user watermarking and present our initial solution to it. More powerful user watermarking techniques are a promising future direction.
@@ 4.1. Problem Definition
We define the problem of user watermarking as to inject a watermark into a target user’s interaction sequence, so that a model trained on the dataset that involves this watermarked user would have a high watermark success rate. We assume that the target user does not have access to other users’ interactions, i.e., the target user cannot view or edit other users’ interactions. The target user is only capable of determining his or her own interaction sequence by interacting with any item at any predetermined time slot. We also assume that the target user does not have any knowledge regarding the recommender system. However, we assume that the target user has access to the item popularity, e.g., how many times an item has been interacted by all users. This is a reasonable assumption on most e-commerce and social platforms (e.g., a user can see how many times an item has been purchased). The target user can leverage this information to help design the watermark.
The user watermark should also meet the universality, unnoticeability, and discriminability requirements mentioned in Sec. 3.1. We do not consider robustness in this work, as we assume that the server would honestly train their model with the raw dataset. Generally, things become more challenging than dataset watermarking as we are only permitted to inject the watermark into one sequence with limited prior knowledge. Next, we present our analysis and solution to this problem.
4.2. DWRS-U
We define the user watermark in a similar way as the dataset watermark. The user watermark is a sequence of watermark items $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l}\}$ with length $l$. $S_{wm}$ would only be inserted into the target user’s interaction sequence.
4.2.1. Watermark Item Selection
Since the watermark signal strength is largely reduced compared with dataset watermark as it only appears once in the entire dataset, we need to compensate for the strength loss by other means, e.g., choices of watermark items, the insert position, as well as the watermark length.
A detailed description of our DWRS-U algorithm for user watermarking is shown in Algorithm 2. Since the watermark only appears once in the entire dataset, it is difficult for the model to memorize it. To alleviate this issue, the watermark items should ideally not appear in other sequences, thereby facilitating the memorization of the watermark pattern. Therefore, we select the most unpopular items as watermark items. Specifically, if the most unpopular item has three interactions, then all those items with three interactions are considered as candidate watermark items. If the number of items with three interactions are smaller than the watermark length $l$, we can expand the candidate item pool to include those items with four interactions. Since users have access to the item popularity, this strategy is feasible to implement.
@@ 4.2.1. Watermark Item Selection
12:  Sample $l$ items from the most unpopular items
@@
4.2.2. Watermark Length & Insert Position
@@
A larger watermark length $l$ naturally enhances memorization, as it provides more dependencies suggested by Fig. 3. This would make the prediction of $i^{wm}_{l}$ easier to overfit given $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$. In fact, we have empirically verified that, under the default setting of our experiment, by setting $l>30$ and using unpopular items as watermark items, the watermark is almost guaranteed to be successful with Recall@10 $>$ 0.9, regardless of the insert position. It suggests that we can always create a strong watermark by increasing the watermark length $l$. However, a long watermark would fail the requirement of unnoticeability and may increase the risk of being detected. Therefore, we next discuss the possibility of choosing a dedicated insert position which can help alleviate the need for a long watermark.
@@
If we want to reduce the watermark length $l$, we need to compensate for the dependency loss by other means. Since we are evaluating the watermark by predicting $y_{wm}=i^{wm}_{l}$ given $x_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$, let us now focus on the last watermark item $y_{wm}=i^{wm}_{l}$. Fig. 3 suggests that items before $i^{wm}_{l}$ contribute significantly to its embedding. However, note that $i^{wm}_{l}$ also contributes to the embeddings of items after it. For convenience, we name a sequence of those items after $i^{wm}_{l}$ as $\mathcal{C}_{n}$, where $n$ denotes the number of items considered. Since the embeddings of items in $\mathcal{C}_{n}$ are obtained by aggregating information from previous items, learning to predict items in $\mathcal{C}_{n}$ in the training process would require updating the embedding of $i^{wm}_{l}$ as well. Therefore, $i^{wm}_{l}$ and $\mathcal{C}_{n}$ would also establish strong correlations, and we may compensate for the dependency loss of reducing $l$ by utilizing $\mathcal{C}_{n}$ to increase the dependency capacity. To achieve this, items in $\mathcal{C}_{n}$ should also be unpopular, as their embeddings are less robust and can be easily fused with the watermark information, thereby increasing the dependency capacity of $S_{wm}$. To this end, we compute the average item popularity of each subsequence, and name the one with the least average popularity as the most unpopular subsequence of the target sequence. We insert the watermark right before the most unpopular subsequence. For example, if given $n$=10, we would first iterate through the target user sequence and extract all subsequences with length 10. Then for each subsequence, we compute the average popularity (number of interactions) of items in it, and select the most unpopular subsequence as $\mathcal{C}_{n}$. The watermark would be inserted right before $\mathcal{C}_{n}$. In this way, we may shorten the length $l$ of the watermark $S_{wm}$. As $l$ decreases, the dependency capacity of $S_{wm}$ also decreases. Thus the model would try to compensate for the capacity loss by utilizing items in $\mathcal{C}_{n}$. When evaluating the watermark, we would discard $\mathcal{C}_{n}$ and append $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{l-1}\}$ to the query sequence and evaluate the prediction of $i^{wm}_{l}$.
@@ 4.2.3. Discussions
Some considerations may be raised regarding the above method. On the one hand, it seems that we may manually insert some unpopular items after $i^{wm}_{l}$ to serve as $\mathcal{C}_{n}$. However, we argue that this does not achieve the goal of reducing the watermark length. Although $S_{wm}$ itself is indeed shortened, we still have to insert $n$ more items to form $\mathcal{C}_{n}$, and the total number of inserted items would become $l+n$. On the other hand, it seems that inserting $S_{wm}$ after the most unpopular subsequence can achieve the same effect of increasing the dependency capacity. But putting unpopular item before the watermark is actually equivalent to expanding the watermark with those items. One still has to include those items in the query when evaluating the watermark in order to achieve a high watermark validity. We are also aware of a weakness of this method. Specifically, if the target user does not have a subsequence filled with unpopular items, or if those items are not unpopular enough, this method may fail and we may have to increase the length of the watermark.
It seems that we could also use DWRS-U for the dataset watermarking. But as we will see in our experiment results, DWRS-D already achieves a very good watermarking quality (high watermark validity and model utility) with a shorter watermark length and a smaller insert ratio. In this case, using DWRS-U to do dataset watermarking would not bring a significant improvement on the watermark validity, but would instead increase the chance of exposure, because the sampling strategies in DWRS-U can introduce potential bias that could be detected. Therefore, DWRS-D is generally preferred for dataset watermarking.
Table 1. Dataset statistics.